A box contains 3 blue cards and 3 white cards. If two cards are drawn one at a time, find the probability that both cards are blue if the draws are made as follows: a) with replacement b) without replacement
Answer by stanbon(75887)

A box contains 3 blue cards and 3 white cards. If two cards are drawn one at a time, find the probability that both cards are blue if the draws are made as follows:
a) with replacement
P(2 blue with replacement) = (3/6)^2 = 1/4
----------------------------
b) without replacement
P(2 blue without replacement) = 3C2/6C2 = 3/15 = 1/5
======================================================
Cheers,
Stan H.